OpenStudy (anonymous):

lx-3l=<2 (Solve for x and sketch the solution set) So confused smh

OpenStudy (anonymous):

plot the graph of y=|x-3| and y=2 . finally shade the region that is below the line y=2 and above the v-graph of |x-3| @ x=3 ps:note i'm telling a v-graph for |ax+b| is at x=-b/a

OpenStudy (anonymous):

-2<=x-3<=2 x>=1 and x<=5 so 1<=x<=5. (|x|<a means -a<x<a

OpenStudy (anonymous):

You can do what brackett suggests also you will get the same thing this is analytical method of solvind mod.